OpenStudy (anonymous):

Solve. 3x2 – 24x = –12

OpenStudy (anonymous):

thanks

myininaya (myininaya):

I'm afraid this a little wrong

myininaya (myininaya):

\[x^2-8x+16=-4+16\] \[(x-4)^2=12\]

OpenStudy (anonymous):

The person above is correct, im afraid i mislead you

OpenStudy (anonymous):

its fine thank u for trying